# Active Label Correction for Building LLM-based Modular AI Systems

### v2

@@ 5 Conclusion
Modular AI agents can leverage LLMs as zero-shot or few-shot learners for rapid deployment without training but there is a large performance gap when compared to fine-tuned models. The proposed pipeline utilizes ALC to improve modular AI agents by leveraging LLM-annotated data and improving its quality to train a replacement model for each module. Compared to synthetic noise models used in previous applications of ALC to the vision domain, we found the noise characteristics of ChatGPT annotated text data to be much more complex. We achieve the same performance as ground truth data using human feedback on a fraction of the dataset. Furthermore, we found this fraction to be less than the fraction of misannotated data for three unique tasks spanning different levels of complexity and data sizes.
6 Limitations
(i) Most publicly available datasets have small text lengths while real applications may involve much longer (conversational) texts. Deploying an AI agent and collecting real datasets is outside the scope of this paper. (ii) Most commercial AI agents allow users to provide feedback through a thumbs-up/down button and this feedback may be accommodated as a signal for misannotation prediction in modular AI agents. We leave this as an open problem for future work. (iii) In real applications, outliers or completely irrelevant examples should be removed from the training dataset, and exploring methods for identifying such examples in LLM-annotated data is another problem that should be addressed in the future work.
@@ Appendix A Extension to Generative Models
$m(x,y_{1:n})=1-P(y_{1:n}|x)$
Generative tasks are far more complex than discriminative tasks. One can expect that the amount of data required to train a model for misannotation prediction for generative tasks will be much more than for discriminative tasks. In future work, we will examine the data size required to effectively perform annotation prediction for different generative tasks.
@@ Appendix B Training Details
Appendix C ChatGPT Prompts for Evaluation Tasks
@@ C.1 ATIS
ATIS (Airline Travel Information System) Intent Classification dataset Hemphill et al. (1990) is labeled with 17 unique intents with one or more intents assigned to user queries. There are only 40 examples with multiple labels. We removed them from the dataset leading to a single-label classification task. We provided ChatGPT with the system message below to instruct it for classification.
@@ C.2 CoNLL
To extract named entities from examples in CoNLL 2003 dataset Tjong et al. (2003), we provided ChatGPT with one example of entity recognition output to describe the output pattern. First, we used the user message below to provide the types of entities that need to be extracted.
@@ C.3 QNLI
QNLI (Question-Answering Natural Language Inference) dataset Wang et al. (2018) has pairs of questions and passages based on the Stanford Question Answering Dataset and the label 00 if the passage answers the question and $1$ otherwise. We use the system message below to prime ChatGPT with the task.
@@
Next, we use the following user message to give the input question and passage and prompt ChatGPT to say ‘YES’ if the passage can answer the question and ‘NO’ otherwise.
@@ Appendix E Noise Models
For random noise, we randomly select examples for the noising process. For each selected example, we change the label to any other label with equal probability. In other words, we assume that each example is equally likely to be mislabeled and each label is equally likely to be replaced by another label.
For label-conditional noise, we first train a classification head on top of the BERT model (‘bert-base-uncased’ on HuggingFace) and use its probability outputs on training examples to calculate label transition probabilities. For each label, we calculate the transition probabilities by averaging the output distribution of examples with that label. After this, we randomly select examples for noising and change their labels based on the transition probabilities. This process assumes that each example is equally likely to be noisy but noisy label is conditionally dependent on the true label. See Li et al. (2022) (Section 3.1) for an introduction to label transition matrix as a representation for label-conditional noise model.

### v3

@@ 5 Conclusion
Modular AI systems can leverage LLMs as zero/few-shot learners for rapid deployment without any training or fine-tuning but there is a large performance gap when compared to fine-tuned models. The proposed pipeline utilizes our method ALC3 to improve modular AI systems by leveraging LLM-annotated data and improving its quality to train a replacement model for each module. Compared to random or label-conditioned synthetic noise models used in previous work, we found the noise characteristics of GPT-3.5-annotated text data to be far more complex. With ALC3, we achieve the same performance as ground truth data using human feedback on a fraction of the dataset. We found this fraction to be 17-24% less than the fraction of misannotated data for three unique tasks spanning different levels of complexity and data sizes, and $\approx 16\%$ better than previous methods.
6 Limitations and Future Work
Most publicly available datasets have small text lengths while real applications may involve much longer (conversational) texts. Deploying an AI system and collecting real datasets is outside the scope of this paper. In future work, we will apply ALC3 to datasets created using data traces obtained from a deployed Jill Watson Kakar et al. (2024) to evaluate its performance after replacing LLM calls with smaller fine-tuned models.
Most commercial conversational AI agents allow users to provide feedback through a thumbs-up/down button and this feedback may be accommodated as a signal for misannotation prediction. We leave this as an open problem for future work.
We used only GPT-3.5 as LLM for annotations. While our results should generalize to different LLMs since this only affects the initial parameters of our experiments (i.e. the noise level of the labelled datasets), more insights could be gained by testing with other LLMs.
@@ Appendix A Extension to Generative Models
$m(x,y_{1:n})=1-P(y_{1:n}|x)^{\frac{1}{n}}$
Since generative tasks are far more complex than discriminative tasks, one can expect that the amount of data required to train a model for misannotation prediction for generative tasks will be much larger than data size required for discriminative tasks. Further, obtaining human feedback or creating a reasonable proxy for human feedback for generative tasks is another difficult challenge. In future work, we will examine the data size required to effectively perform annotation prediction for different generative tasks.
In previous work in human-in-the-loop learning for generative tasks, Reinforcement Learning with Human Feedback (RLHF) used for ChatGPT OpenAI (2022) provides an answer to a part of the problem but doesn’t address how to select useful examples where human-feedback will be most helpful in terms of improving model performance.
@@ Appendix B Training Details
Appendix C GPT-3.5 Prompts for Evaluation Tasks
@@ C.1 ATIS
ATIS (Airline Travel Information System) Intent Classification dataset Hemphill et al. (1990) is labeled with 17 unique intents with one or more intents assigned to user queries. There are only 40 examples with multiple labels. We removed them from the dataset leading to a single-label classification task. We provided GPT-3.5 with the system message below to instruct it for classification.
@@ C.2 CoNLL
To extract named entities from examples in CoNLL 2003 dataset Tjong et al. (2003), we provided GPT-3.5 with one example of entity recognition output to describe the output pattern. First, we used the user message below to provide the types of entities that need to be extracted.
@@ C.3 QNLI
QNLI (Question-Answering Natural Language Inference) dataset Wang et al. (2018) has pairs of questions and passages based on the Stanford Question Answering Dataset and the label 00 if the passage answers the question and $1$ otherwise. We use the system message below to prime GPT-3.5 with the task.
@@
Next, we use the following user message to give the input question and passage and prompt GPT-3.5 to say ‘YES’ if the passage can answer the question and ‘NO’ otherwise.
@@ Appendix E Noise Models
For random noise, we randomly select examples for the noising process. For each selected example, we change the label to any other label with equal probability. In other words, we assume that each example is equally likely to be mislabeled and each label is equally likely to be replaced by another label. Kremer et al. (2018) uses this process to add noise to clean datasets for experimentation.
For label-conditional noise, we first train a classification head on top of the BERT model (‘bert-base-uncased’ on HuggingFace) and use its probability outputs on training examples to calculate label transition probabilities. For each label, we calculate the transition probabilities by averaging the output distribution of examples with that label. After this, we randomly select examples for noising and change their labels based on the transition probabilities. This process assumes that each example is equally likely to be noisy but noisy label is conditionally dependent on the true label. Li et al. (2022) use the same process for introducing noise into clean datasets in their experiments.